OpenStudy (anonymous):

Kevin invests $100 in an account that triples every 2 years. Anna invests $1000 in an account that doubles every 3 years. How many years will their accounts be equal in value? Assuming that no deposits or withdrawals were made. Anyone? Please please please help me :) :3

OpenStudy (anonymous):

Try writing both of them as a function, then find the point where they intercept.

OpenStudy (anonymous):

Let's look at Kevin first

OpenStudy (anonymous):

He starts with 100, then 2 years later he has 300, and 2 years later he has 900 (the 300 gets tripled, not the 100)

OpenStudy (anonymous):

So can we find a function that has such behavior?

OpenStudy (anonymous):

@zimmah but there's a formula for this right? Is it effective to use formulas such has, A = P(1 + r/n)^nt and A=Pe^rt ??

OpenStudy (anonymous):

ohhh wait, i see, so it's not the 100 that gets tripled.

OpenStudy (anonymous):

Yes it is for the first 2 years.

OpenStudy (anonymous):

But then, once you have 300, the 300 will get tripled 2 years later.

OpenStudy (anonymous):

You can use a simular equation

OpenStudy (anonymous):

To the one you typed.

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

So how will i put those into an equation? :)

OpenStudy (anonymous):

I have to equate both of the equations to find t right? t = time :)

OpenStudy (anonymous):

100(1+2)^(t/2) I believe, try it and see if it gives the values we would expect.

OpenStudy (anonymous):

@zimmah you used A = P(1+r/n)^nt ?? :)

OpenStudy (anonymous):

Yes, check if the one I created is correct, the try to make one for Anna, than if you have both you ca type an = sign bet gg them and solve for 0

OpenStudy (anonymous):

@zimmah what confuses me is the r. What's the r? Haha.

OpenStudy (anonymous):

do we use r as constant?

OpenStudy (anonymous):

The r is the rate in % of increase.

OpenStudy (anonymous):

ahhh okay :) so r is 3 for kevin?

OpenStudy (anonymous):

wait, i still dont get it. haha pardon meee

OpenStudy (anonymous):

Well, not exactly

OpenStudy (anonymous):

oh coz r is %?

OpenStudy (anonymous):

Let me start my computer any I will expai

OpenStudy (anonymous):

okay i'll wait :)

OpenStudy (anonymous):

ok, here we go

OpenStudy (anonymous):

looking at the increase from year 0 to year 2, kevin has 100, and it increase from 100 to 300.

OpenStudy (anonymous):

that's a 300% increase.

OpenStudy (anonymous):

so if you write it in the form 100(1+r) than r would be 2, because 1+2=3 (note that it are percentages, and 3 actually means 300%

OpenStudy (anonymous):

you understand thus far?

OpenStudy (anonymous):

Okaay i do :)

OpenStudy (anonymous):

alright, now notch it up a step.

OpenStudy (anonymous):

we can't simply say that after 4 years (2 periods of 2 years) his money will increase with 600%. can you see why?

OpenStudy (anonymous):

because it's not the 100 that triples :)

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

it's the 300 :)

OpenStudy (anonymous):

so we have to make an equation that makes a 'copy' of itself, that sort of 'remembers' the value it has in between two years and than uses a copy of itself with the new value and repeats it for n amount of times where n is the amount of periods you need.

OpenStudy (anonymous):

this may sound hard, but once i show you, it will probably be clear.

OpenStudy (anonymous):

\[100(1+2)=300\] \[300(1+2)=900\] 900(1+2)=2700

OpenStudy (anonymous):

you see the (1+2) is the thing that always stays constant, and it what is causing the tripling.

OpenStudy (anonymous):

so we want to 'copy' that part of the formula n amount of times. (where n is the amount of periods we want to know)

OpenStudy (anonymous):

what's the mathematical symbol for making n amount of copies of itself as factors?

OpenStudy (anonymous):

if i want for example 3 copies of 2 that multiply with each other, how could i write it?

OpenStudy (anonymous):

so 2*2*2, how can i write it mathematically?

OpenStudy (anonymous):

2^3??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, since every two years our the percentage part of our equation needs to make a copy of itself, we can use ^(t/2) as our factor.

OpenStudy (anonymous):

\[\Large \left( 1+2 \right)^{\frac{ t }{ 2 }}\] works like a copying machine, it will print you t/2 times (1+2)

OpenStudy (anonymous):

and we do the same for Anna?

OpenStudy (anonymous):

yes, try it yourself, so we can see if you can put what you learned to practise.

OpenStudy (anonymous):

\[100(3)^{t/2} = 1000(2)^{t/3}\] uhmmm, like this? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so im gonna use natural log to solve it right? :D

OpenStudy (anonymous):

and use my scientific calculator haha..

OpenStudy (anonymous):

uhm, i got 7.23 years.

OpenStudy (anonymous):

\[\large 100(3)^{\frac{ t }{ 2 }}-1000(2)^{\frac{ t }{ 3 }}=0\]

OpenStudy (anonymous):

or you can use that indeed lol

OpenStudy (anonymous):

i got the same

OpenStudy (anonymous):

@zimmah Thank you!! :)

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

am i right or what but i can see from the equation that it has a relation to exponential growth formula :) @zimmah

OpenStudy (anonymous):

yes, it does.

OpenStudy (anonymous):

it's a variation of it.

OpenStudy (anonymous):

ohhhhh, i see :)

OpenStudy (anonymous):

well thanks for making me realize everything LOL

OpenStudy (anonymous):

it's ok, i like it when people truly understand what is going on instead of just copying formulas without knowing what the formula does, people enjoy it a lot more when they know what is going on.

OpenStudy (anonymous):

i highly agree with you, sir! haha :)